XYZ Co. operates in a competitive market. Its Total Product (Q) is given as Q=f(L)= 2L, and it

takes the wage and price as given. Derive the firm's short-run demand for labor as a function of w

and p. How much labor will the firm hire if W=₵20 and P=₵100?


1
Expert's answer
2021-11-07T19:51:48-0500

MPL = W / p = w

MPL="\frac{w}{p}" =w

MRPL = w

MPL "\times" P = w

MPL = 2L

P = 100

W = 20

3L  "\times" 100 = 20

300L = 20

L = 0.0667

Q = 2L = 2 "\times" 0.0667 = 0.133

Labor = 0.133

The firm will hire 0.133 labor